OpenStudy (anonymous):

1.what is the value of K if the line 3x-2y=1 is perpendicular to the line y-(2K+1)x=3?

OpenStudy (shubhamsrg):

product of slopes of perpendicular lines = -1

OpenStudy (anonymous):

yup

OpenStudy (shubhamsrg):

So where are you stuck ?

OpenStudy (anonymous):

uhm.. EVERYTHING??? ahahaha.xD

OpenStudy (shubhamsrg):

Can you tell me the slope of the line 3x-2y = 1 ?

OpenStudy (shubhamsrg):

Hint : try to compare with y= mx+c form, where m is the slope

OpenStudy (anonymous):

wait,,

OpenStudy (anonymous):

negative 1/y???

OpenStudy (shubhamsrg):

m should be a constant. try again.

OpenStudy (anonymous):

,, aw, my brain hurts,

OpenStudy (shubhamsrg):

Sure it does. hmm

OpenStudy (shubhamsrg):

Leme guide you through 1st line, you try on your own for second.

OpenStudy (anonymous):

okaaay, thanks,

OpenStudy (shubhamsrg):

3x - 2y =1 3x = 2y+1 (3/2)x = y + (1/2) y = (3/2)x - (1/2) Did you follow ?

OpenStudy (nurali):

y-(2K+1)x=3 adding (2K+1)x to both sides we get y = (2K+1)x +3

OpenStudy (anonymous):

so the slope is -3

OpenStudy (shubhamsrg):

from where did you pick -3 ?

OpenStudy (nurali):

y = 3/2x-1/2 what is the slope now??

OpenStudy (anonymous):

oh.xD hahaha the slope is 3/2

OpenStudy (nurali):

yes exactly

OpenStudy (nurali):

y = (2K+1)x +3 and the here the slope is ??

OpenStudy (nurali):

reply.

OpenStudy (anonymous):

waiiit,,xD

OpenStudy (anonymous):

2k??/

OpenStudy (nurali):

no

OpenStudy (nurali):

try it again

OpenStudy (anonymous):

waiit,, isnt the slope for the 1st eqAuation 3??

OpenStudy (anonymous):

-2y=-3x+1 (-2y=-3x+1)-1 2y=3x-1 so slope is 3

OpenStudy (anonymous):

oh, you're gonna divide it by 2

OpenStudy (nurali):

yes

OpenStudy (anonymous):

the second equation is sooo confusing.

OpenStudy (nurali):

y = (2K+1)x +3 and the here the slope is ?? formula is y=mx+b here m = slope and b = y-intercept

OpenStudy (anonymous):

(2K+1)????

OpenStudy (nurali):

yes exactly

OpenStudy (anonymous):

i thought, you were, going to multiply it from the x . hahaha

OpenStudy (anonymous):

then how to get the value of k??

OpenStudy (nurali):

m1m2=-1 now these line are perpendicular to each other thus their slopes are related as m1 = -1/m2 here m1 = 3/2 and m2 = 2k+1 3/2 = -1/(2k+1) can you solve for K now???

OpenStudy (anonymous):

-5/6???

OpenStudy (nurali):

sorry it correct so K = -5/6 is the required value of K

OpenStudy (anonymous):

okay,,xD

OpenStudy (anonymous):

.oh,, so i'm finished,, thaaaaaaanks!!!!!!!

OpenStudy (nurali):

My Pleasure.